OpenStudy (anonymous):

I've got part A, can someone help me out? -16t^2+50t+90 is the equation and answer to Part A. Part B: What is the maximum height that the projectile will reach? Show your work.

OpenStudy (cwrw238):

have you done any calculus finding derivatives?

OpenStudy (anonymous):

a little bit i guess

OpenStudy (cwrw238):

you need to differentiate the above function and equate it to zero then solve for t

OpenStudy (anonymous):

i got 1.5625

OpenStudy (cwrw238):

this will give you the value of t when projectile is at maximum height you then plug this into the original function of t

OpenStudy (cwrw238):

-32t + 50 = 0 t = 50/32 = 1.5625 now plug this into -16t^2 + 50t + 90

OpenStudy (anonymous):

-16(1.5625)^2+50(1.5625)+90 = 129.0625 would that be correct?

OpenStudy (cwrw238):

129.06?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

okay thank you! :)